Lily Tucker (single) owns and operates a bike shop as a sole proprietorship. In 2019, she sells the following long-term assetLily Tucker (single) owns and operates a bike shop as a sole proprietorship. In 2019, she sells the following long-term assets used in her business: Asset Sales Price Cost Accumulated Depreciation Building $234,200 $204,200 $56,200 Equipment 84,200 152,200 27,200 Lily's taxable income before these transactions is $194,700. What are Lily's taxable income and tax liability for the year? Use Tax Rate Schedule for reference. (Do not round intermediate calculations. Round your answers to the nearest whole dollar amount.)

2019 Tax Rate Schedules

Individuals

Schedule X-Single

If taxable income is over: But not over: The tax is:
$           0 $    9,700 10% of taxable income
$    9,700 $ 39,475 $970 plus 12% of the excess over $9,700
$ 39,475 $ 84,200 $4,543 plus 22% of the excess over $39,475
$ 84,200 $160,725 $14,382.50 plus 24% of the excess over $84,200
$160,725 $204,100 $32,748.50 plus 32% of the excess over $160,725
$204,100 $510,300 $46,628.50 plus 35% of the excess over $204,100
$510,300 $153,798.50 plus 37% of the excess over $510,300

Schedule Y-1-Married Filing Jointly or Qualifying Widow(er)

If taxable income is over: But not over: The tax is:
$           0 $ 19,400 10% of taxable income
$ 19,400 $ 78,950 $1,940 plus 12% of the excess over $19,400
$ 78,950 $168,400 $9,086 plus 22% of the excess over $78,950
$168,400 $321,450 $28,765 plus 24% of the excess over $168,400
$321,450 $408,200 $65,497 plus 32% of the excess over $321,450
$408,200 $612,350 $93,257 plus 35% of the excess over $408,200
$612,350 $164,709.50 plus 37% of the excess over $612,350

Schedule Z-Head of Household

If taxable income is over: But not over: The tax is:
$           0 $ 13,850 10% of taxable income
$ 13,850 $ 52,850 $1,385 plus 12% of the excess over $13,850
$ 52,850 $ 84,200 $6,065 plus 22% of the excess over $52,850
$ 84,200 $160,700 $12,962 plus 24% of the excess over $84,200
$160,700 $204,100 $31,322 plus 32% of the excess over $160,700
$204,100 $510,300 $45,210 plus 35% of the excess over $204,100
$510,300 $152,380 plus 37% of the excess over $510,300

Schedule Y-2-Married Filing Separately

If taxable income is over: But not over: The tax is:
$           0 $    9,700 10% of taxable income
$    9,700 $ 39,475 $970 plus 12% of the excess over $9,700
$ 39,475 $ 84,200 $4,543 plus 22% of the excess over $39,475
$ 84,200 $160,725 $14,382.50 plus 24% of the excess over $84,200
$160,725 $204,100 $32,748.50 plus 32% of the excess over $160,725
$204,100 $306,175 $46,628.50 plus 35% of the excess over $204,100
$306,175 $82,354.75 plus 37% of the excess over $306,175
Answer #1

Answer :-

Selling Price (f) Cost (a) Accumulated Depreciation(b) Book Value(d) =(a)-(b) Profit(loss)(g)=(f)-(d)
Building                 234,200          204,200                     56,200           148,000           86,200
Equipment                   84,200          152,200                     27,200           125,000         (40,800)
Total Profit(loss)           45,400

Taxable income excluding capital assets loss or profit = $194,700

Profit for sale of capital assets = $45,400

Taxable Income = 194,700+45,400 = $240,100

Calculation of Tax :-

Lily is single individual so we have to calculate tax as per individual Schedule X -Single Table

If taxable income is over: But not over: The tax is:
$           0 $    9,700 10% of taxable income
$    9,700 $39,475 $970 plus 12% of the excess over $9,700
$39,475 $84,200 $4,543 plus 22% of the excess over $39,475
$84,200 $160,725 $14,382.50 plus 24% of the excess over $84,200
$160,725 $204,100 $32,748.50 plus 32% of the excess over $160,725
$204,100 $510,300 $46,628.50 plus 35% of the excess over $204,100
$510,300 $153,798.50 plus 37% of the excess over $510,300

Tax on income $240,100 = 46,628.50 + (240,100-204,100)*35%

Tax on income $240,100 = 46,628.50 + 12,600

Tax on income $240,100 = $59,228.50 or $59,229(rounded off)